OpenStudy (adamaero):

If x->infinity, can L'hopital's rule be applied: sin(x)/(1+sqrt(x)) ?

OpenStudy (adamaero):

It's infinity over infinity right now?

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

limits are not my strong point i'm afraid but l'hopitals seems the wat to go

OpenStudy (rsadhvika):

sinx is bounded between -1 and 1, can u check again if its really in infinity/infinity form ?

OpenStudy (cwrw238):

yes thats true - its not in infinity/ infinity form since the denominator is x^ (1/2) the wouldn't the limit be 0?

OpenStudy (cwrw238):

yes 0 is the answer

OpenStudy (cwrw238):

as rsadyvica said the values of sine lies between -1 and 1 and as x approaches infinity so does sqrt x

OpenStudy (redheadangel):

if you need extra help I recomend looking at the Khan academy video for this, he does a whole vid on the exact thing you need help with :)

OpenStudy (adamaero):

ya, you guys were right--0--but I was suppose to use squeeze theorem

OpenStudy (rsadhvika):

\[\large -1\le\sin x \le 1 \implies \dfrac{-1}{1+\sqrt{x}}\le \dfrac{\sin x}{1+\sqrt{x}}\le \dfrac{1}{1+\sqrt{x}}\]

OpenStudy (rsadhvika):

take the limit through out

OpenStudy (rsadhvika):

show that the limit in question is bounded between 0 and 0 that proves that the limit is exactly 0